OpenStudy (anonymous):

ASAP HELP ME PLEASE! What is the distance between (6, 2) and (-3, -2)? 9 3 sqrt97 5

OpenStudy (anonymous):

\[distance = \sqrt{(x2 - x1)^2 + (y2 - y1)^2}\]

OpenStudy (anonymous):

\[\sqrt{((-3) - 6)^2 + ((-2) - 2)^2}\] \[\sqrt{(-9)^2 + (-2)^2}\] \[\sqrt{81 + 16}\] \[\sqrt{97}\]